# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws (Extended Version)

### v2

@@ VII Numerical Experiments
VII-A Representating a Feasible BBS Trajectory
@@
The constraints are $\left|u\right|\leq 1$, $x_{1}\geq-0.7$, and $x_{3}+x_{4}\leq 0.5$. As defined in Proposition 2, $x_{1}\geq-0.7$ is a 3rd-order state constraint. As shown in Figs. 4(a-c), a feasible BBS trajectory ${\boldsymbol{x}}_{\text{ASL}}$ consists of full features defined in this paper. The ASL is ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}{\mathcal{S}}_{4}\left({\mathcal{S}}_{5},\left\{{\mathcal{T}}_{5}^{(1)}\right\}\right){\mathcal{S}}_{6}{\mathcal{E}}_{6}{\mathcal{S}}_{7}{\mathcal{S}}_{8}$. Specifically, ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{4}$, and ${\mathcal{S}}_{8}$ are unconstrained arcs where $u\equiv-1$. ${\mathcal{S}}_{2}$, ${\mathcal{S}}_{5}$, and ${\mathcal{S}}_{7}$ are unconstrained arcs where $u\equiv+1$. ${\mathcal{S}}_{3}$ and ${\mathcal{S}}_{6}$ are constrained arcs where $x_{3}+x_{4}\equiv 0.5$. ${\mathcal{T}}_{5}^{(1)}$ is a tangent markers where $x_{1}$ is tangent to $-0.7$. ${\mathcal{E}}_{6}$ is an additional end-constraint requiring $u\left(t_{6}^{-}\right)=1$. In other words, the control $u$ is continuous at the connection of the constrained arc ${\mathcal{S}}_{6}$ and the unconstrained arc ${\mathcal{S}}_{7}$. In contrast, $u$ is not continuous at the two ends of another constrained arc ${\mathcal{S}}_{3}$. One can check that Theorem 3 holds for ${\boldsymbol{x}}_{\text{ASL}}$.
@@
VII-B Determining the Optimality of a 4th-Order Trajectory
A near-optimal trajectory planning method is proposed in our previous work [20]. Consider a 4th-order trajectory for high-order chain-of-integrator systems (48a) provided in [20] where ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, and ${u_{\mathrm{m}}}=1$, as shown in Fig. 5(b). The ASL of the original trajectory ${\boldsymbol{x}}\left(t\right)$ is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$. Note that $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$; hence, Corollary 2 implies that the terminal time ${t_{\mathrm{f}}}=9.8604$ is not optimal. Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$. Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns has full row rank. Then, perturb $t_{4}$ to $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (48a). As shown in Fig. 5(a), $t_{\mathrm{f}}^{\prime}$ achieves its local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$. It can be observed in Fig. 5(b2) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory. Hence, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint. Specifically, ${\boldsymbol{x}}^{\prime}$ achieves a 0.06% reduction in the terminal time compared to ${\boldsymbol{x}}$. Therefore, the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information.
@@ VII-C Optimizing a 5th-Order Feasible BBS Trajectory
Figure 7: Comparison of the original trajectory ${\boldsymbol{x}}_{1}$ [36], the optimized trajectory ${\boldsymbol{x}}_{6}$ through the proposed state-centric necessary condition, and the trajectory ${\boldsymbol{x}}_{\text{base}}$ planned by [37]. (a) The states and controls. (b) The terminal time. (c) The total variation of the control. (d) The open-loop error of terminal states.
VIII Conclusions
This paper has set out to establish an innovative theoretical framework for time-optimal control of controllable linear systems with a single input. The proposed augmented switching law (ASL) represents the input control and the feasibility of a bang-bang and singular (BBS) trajectory in a compact form. Specifically, the equality and inequality constraints induced by the ASL represent a sufficient and necessary condition for the trajectory’ feasibility. Based on the ASL, a given feasible trajectory can be perturbed with feasibility guarantees, resulting in a state-centric necessary condition for time optimality since any perturbed feasible trajectory should have a strictly longer terminal time than the optimal one. The proposed necessary condition states that the Jacobian matrix induced by the ASL must not be full row rank. In contrast to traditional costate-based conditions, the developed necessary condition requires only states without dependence on costate information.

### v3

@@ VII Numerical Experiments
VII-A Representing a Feasible BBS Trajectory
@@
The constraints are $\left|u\right|\leq 1$, $x_{1}\geq-0.7$, and $x_{3}+x_{4}\leq 0.5$. As defined in Proposition 2, $x_{1}\geq-0.7$ is a 3rd-order state constraint. As shown in Figs. 4(a-c), a feasible BBS trajectory ${\boldsymbol{x}}_{\text{ASL}}$ consists of full features defined in this paper. The ASL is ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}{\mathcal{S}}_{4}\left({\mathcal{S}}_{5},\left\{{\mathcal{T}}_{5}^{(1)}\right\}\right){\mathcal{S}}_{6}{\mathcal{E}}_{6}{\mathcal{S}}_{7}{\mathcal{S}}_{8}$. Specifically, ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{4}$, and ${\mathcal{S}}_{8}$ are unconstrained arcs where $u\equiv-1$. ${\mathcal{S}}_{2}$, ${\mathcal{S}}_{5}$, and ${\mathcal{S}}_{7}$ are unconstrained arcs where $u\equiv+1$. ${\mathcal{S}}_{3}$ and ${\mathcal{S}}_{6}$ are constrained arcs where $x_{3}+x_{4}\equiv 0.5$. ${\mathcal{T}}_{5}^{(1)}$ is a tangent marker where $x_{1}$ is tangent to 0.7-0.7. ${\mathcal{E}}_{6}$ is an additional end-constraint requiring $u\left(t_{6}^{-}\right)=1$. In other words, the control $u$ is continuous at the connection of the constrained arc ${\mathcal{S}}_{6}$ and the unconstrained arc ${\mathcal{S}}_{7}$. In contrast, $u$ is not continuous at the two ends of another constrained arc ${\mathcal{S}}_{3}$. One can check that Theorem 3 holds for ${\boldsymbol{x}}_{\text{ASL}}$.
@@
VII-B Determining the Optimality of a 4th-Order Trajectory
A near-optimal trajectory planning method is proposed in our previous work [20]. Consider a 4th-order trajectory for high-order chain-of-integrator systems (26a) provided in [20] where ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, and ${u_{\mathrm{m}}}=1$, as shown in Fig. 5(b). The ASL of the original trajectory ${\boldsymbol{x}}\left(t\right)$ is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$. Note that $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$; hence, Corollary 2 implies that the terminal time ${t_{\mathrm{f}}}=9.8604$ is not optimal. Let $\boldsymbol{H}$} be the equality constraints induced by $\mathcal{L}$. Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns has full row rank. Then, perturb $t_{4}$ to $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (26a). As shown in Fig. 5(a), $t_{\mathrm{f}}^{\prime}$ achieves its local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory $\boldsymbol{x}$$^{\prime}$. It can be observed in Fig. 5(b2) that $x_{3}^{\prime}$ is tangent to $x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory. Hence, the original trajectory fails to fully utilize the feasible set, and the extreme of $t_{\mathrm{f}}$$^{\prime}$ activates a new constraint. Specifically, $\boldsymbol{x}$$^{\prime}$ achieves a 0.06% reduction in the terminal time compared to $\boldsymbol{x}$}. Therefore, the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information.
@@ VII-C Optimizing a 5th-Order Feasible BBS Trajectory
VIII Conclusion
This paper has set out to establish an innovative theoretical framework for time-optimal control of controllable linear systems with a single input. The proposed augmented switching law (ASL) represents the input control and the feasibility of a bang-bang and singular (BBS) trajectory in a compact form. Specifically, the equality and inequality constraints induced by the ASL represent a sufficient and necessary condition for the trajectory’ feasibility. Based on the ASL, a given feasible trajectory can be perturbed with feasibility guarantees, resulting in a state-centric necessary condition for time optimality since any perturbed feasible trajectory should have a strictly longer terminal time than the optimal one. The proposed necessary condition states that the Jacobian matrix induced by the ASL must not be of full row rank. In contrast to traditional costate-based conditions, the developed necessary condition requires only states without dependence on costate information.
@@ VIII Conclusion
Figure 7: Comparison of the original trajectory ${\boldsymbol{x}}_{1}$ [35], the optimized trajectory ${\boldsymbol{x}}_{6}$ based-on the proposed state-centric necessary condition, and the trajectory ${\boldsymbol{x}}_{\text{base}}$ planned by [36]. (a) The states and controls. (b) The terminal time. (c) The total variation of the control. (d) The open-loop error of terminal states.